# New perspectives on the wave equation for a massive scalar field in Schwarzschild spacetime

### v1

@@ VII Large-$r$ solution of the full wave equation
So far, we have studied the behaviour of solutions for fixed frequencies. Since one of our goals is to study the spectrum of a massive field far away from the black hole, we now focus the attention on the study of the solution at spacelike infinity by including the contribution of all frequencies (neglecting the angular part). In the previous sections we have discovered that, when $r\rightarrow\infty$, the Fourier modes take the approximate form (49). The large-$r$ form of the scalar field $\psi$ which solves Eq. (1) and satisfies the asymptotic condition
@@
is therefore expressed by Eq. (4), where, upon defining
@@
The integral is qualitatively different in the $\omega^{2}<\mu^{2}$ case, since the last two terms in Eq. (51) become imaginary, leading to
@@
The investigation of these two integrals is a hard task. However, we can rely on analytical approximations and numerical analysis in order to extract the essential features from them. The starting point is given by studying the behaviour in frequencies of the two integrand functions, which we label as
$\Omega_{l\mu}^{I}(t,r,\omega)=\frac{1}{\sqrt{2\pi r(r-2M)}}e^{-i\varphi(t,r,\omega)}\left(1+\sum_{s=1}^{\infty}\frac{B_{s}(l,\omega\mu)}{r^{s}}\right),$ (54)
@@
Notice that the $l$-dependence is encoded in the coefficients of the series expansion, requiring to add also the angular part (given by spherical harmonics) to study the superposition of several waves. We therefore limit ourselves to the study of the $s$-wave, given by the constraint $B_{s}(l,\omega,\mu)=B_{s}(0,\omega,\mu)$. Moreover, for numerical reasons, we consider only the first five negative powers of $r$, arriving at a semplified version of Eqs. (54), (55). Upon setting
@@
we first consider the plots of the real (Fig. 7) and imaginary (Fig. 8) part of Eq. (55).
Figure 7: Plot of Re$\left[\Omega_{0\mu}^{II}(1,r,\omega)\right]$ for $r\in[100,120]$, $\omega\in[-\mu,\mu]$, setting $\mu=0.1$ and $M=10$, up to the $r^{-5}$ term. Along the $\omega$-axis the function has a Gaussian-like behaviour (it is possible to prove it by expanding in series Eq. 55 for small frequencies), while it diverges along the $r$-axis. The integral, which is of course the area under the curve, is clearly divergent in the $r\rightarrow\infty$ limit.
Figure 8: Plot of Im$\left[\Omega_{0\mu}^{II}(1,r,\omega)\right]$ for $r\in[100,120]$, $\omega\in[-\mu,\mu]$, setting $\mu=0.1$ and $M=10$, up to the $r^{-5}$ term. Unlike the real part, the phase shows a different behaviour by virtue of its odd-parity with respect to $\omega$. The contributions to the integral on the $r$-axis seem to vanish in light of this feature.
It is possible to evaluate numerically the integral given by Eq. (53) by expanding $\Omega_{l\mu}^{II}(t,r,\omega)$ around $\omega=0$, since this region gives the largest contribution to it (as shown in Fig. 7). We have performed such evaluation up to the sixth power of $\omega$ for the choosen parameters, arriving at the conclusion that such a piece diverges. In order to obtain also an analytical estimate of such a divergence, we consider the change of variable given by
@@
applied to the leading term of Eq. (55), obtaining therefore
@@
Even though it is possible to perform more accurate computations (including more terms and/or the negative powers of $r$ of the series expansion) it is clear that the dominant contribution is the one of the exponential term, which causes the divergence of the whole integral in the $r\rightarrow\infty$ limit, as shown by the purely numerical analysis given before. Notice that the field mass term, $\mu$, plays a central role: in the massless case the equivalent of such a piece vanishes (as it happens here in the $\mu\rightarrow 0$ limit). We have therefore to face a case where a field not only does not vanish at spacelike infinity, but furthermore it diverges exponentially. As far as we can see, this property is compatible with the findings in Ref. Huh , where part (II) of Proposition 2.1 therein shows that the norm of a massive scalar field obeying the wave equation in Schwarzschild spacetime is majorized by the product of two functions, one of which is indeed divergent in a suitable limit. This agreement merits explicit mention, because the work in Ref. Huh has exploited advanced methods from the modern theory of hyperbolic equations, whereas we have used the standard techniques of classical mathematical physics.
@@
The case of $\Omega_{l\mu}^{I}(t,r,\omega)$ is strongly different. Upon relying on the same conventions, we show the plots of its real part (the ones for the phase are very similar) for $\omega\in[-10^{3},-\mu]$ in Fig. 9 and for $\omega\in[\mu,10^{3}]$ in Fig. 10.
Figure 9: Plot of Re$\left[\Omega_{0\mu}^{I}(1,r,\omega)\right]$ for $r\in[1,1.2]\cdot 10^{3}$, $\omega\in[-10^{3},-\mu]$, setting $\mu=0.1$ and $M=10$, up to the $r^{-5}$ term. For every fixed value of $r$, the field exhibits increasing oscillations as $\omega$ grows (in this specific case for $\omega<-400$. In the low-frequency limit, the field is neglibile. Both behaviours suggest that the numerical interval is not divergent.
Figure 10: Plot of Re$\left[\Omega_{0\mu}^{I}(1,r,\omega)\right]$ for $r\in[1,1.2]\cdot 10^{3}$, $\omega\in[\mu,10^{3}]$, setting $\mu=0.1$ and $M=10$, up to the $r^{-5}$ term. The discussion of such a figure follows closely the one of Fig. 9.
@@
On both numerical and analytical ground we can say that, while the function $\Omega_{l}^{I}(t,r,\omega)$ diverges, the associated integral (given by Eq. (52)) remains finite and eventually vanishes at spacelike infinity. In fact, we have an oscillating term $e^{-i\varphi}$ alleviated by an infinite series of negative powers of $r$.
@@
Interestingly, we find that a massive scalar field contains a term which diverges at spacelike infinity, while this feature is absent in the massless case. In conclusion, the mass plays a crucial role in the asymptotic description of a scalar field in spherical coordinates.
VIII Application of our Sec. VI to the Zerilli equation
@@
The systematic algorithm of Sec. VI can be applied to equations of even greater interest. For this purpose, we here consider the recent analysis in Ref. Hamaide of the Zerilli equation Zerilli , which is obtained by perturbing a background Schwarzschild metric with a small parity-even perturbation expressed by a symmetric tensor field of type $(0,2)$. Eventually, the work in Ref. Hamaide obtains an inhomogeneous radial wave equation, that we express in the form
@@
where $\eta=\frac{(l-1)(l+2)}{2}$. The work in Ref. Hamaide uses a Green-function method in order to solve the inhomogeneous differential equation (61) by using the well-established technique for obtaining the Green function of the operator on the left-hand side of Eq. (61) from solutions of the associated homogeneus equation
@@
It has been our choice to focus on this particular aspect of the general problem, since we have already written several sections on the massive scalar case, which is the main topic of our paper.
By repeating the reduction to normal form as we did in Sec. II, we obtain $U(r,l,\omega)$ in the factorized form
@@
having defined
@@
At this stage, in close analogy with Eq. (41), we may expand Eq. (68) as $r\rightarrow\infty$ in the form
@@
The formulae for expressing $\zeta$ and $B_{s}$ in terms of the $A_{k}(l,\omega)$ coincide with the ones in Sec. VI, whereas the $A_{k}$ are different. Indeed, we find for example that
@@
Hereafter, we plot the real (Fig. 11) and imaginary (Fig. 12) part of Eq. (66) upon considering the parameters
@@
up to the $B_{5}$ coefficient in Eq. (71). Our findings in Figs. 11 and 12 are in full agreement with the plane-wave behaviour as $r\rightarrow\infty$ obtained in Ref. Hamaide .
@@
Figure 11: The real part of Eq. (66) is shown for $r\in[100,300]$, given the parameters (78) and summing over $s$ from 1 to 5 in Eq. (71). After achieving high values in the region closest to the event horizon, the field shrinks at $r\approx 140$ and then behaves like a plane wave, as it should be.
@@
Figure 12: The imaginary part of Eq. (66) is shown for $r\in[100,300]$, given the parameters (78) and summing over $s$ from 1 to 5 in Eq. (71). Its behaviour is very similar to the one of the real part, leading to the same conclusions.
IX Concluding remarks and open problems
@@
The scalar wave equation in Schwarzschild spacetime has been studied intensively over more than half a century Persides1 ; Persides2 ; Persides3 ; Schmidt ; Fiziev ; Philipp ; Kehrberger , until the very recent and highly valuable work in Ref. Bianchi , but the material in our Secs. VI and VII remains original. More precisely, as far as we know, our original contributions are as follows: (i) For a massive scalar field in Schwarzschild geometry, our Eqs. (47)-(49) and (108)-(109) yield the full Poincaré asymptotic expansion of Fourier modes that solve the radial wave equation. (ii) Our Eqs. (52)-(53) yield the integral representation of the massive scalar field, with the resulting numerical analysis discussed therein.
@@
For massive scalar fields in Schwarzschild spacetime, the literature has mainly focused on late-time behaviour Burko of Fourier modes, whereas rigorous results on solutions of the Cauchy problem are along the lines of Ref. Huh . We think that the next step will consist in applying analogous techniques to other spherically symmetric spacetimes, both in black-hole theory Minucci and in cosmological backgrounds BE . For this purpose, we have eventually studied the homogeneous equation associated to the inhomogeneous Zerilli equation in a Schwarzschild background. Our original formulae (72)-(77) can be applied to build the Green function, which in turn leads to the solution of the inhomogeneous Zerilli equation as shown in Ref. Hamaide .
Acknowledgements
G.E. is grateful to INDAM for membership.
@@ Appendix A The Confluent Heun equation and its solutions
Given a differentiable function $Y$ of the variable $z$, the confluent Heun equation Ronveaux is a second-order ordinary differential equation of the following form:
@@
while $\alpha,\beta,\gamma,\delta,\eta$ are free (in general complex) parameters. A Heun equation has four singular points, while its confluent form has two regular singular points at $z=0$, $z=1$ of $s$-rank $1$ and an irregular singular point at $z=\infty$ of $s$-rank $2$. Indeed, the confluent form is obtained from the standard one by glueing two singularities. In the literature, only power-series solutions in the neighbourhood of singular points are known, resulting in three sets of two linearly independent solutions. The ones in the neighbourhood of $z=0$ and $z=1$ are called Frobenius solutions, while the ones in the neighbourhood of the point at infinity are known as Thomé solutions. Upon denoting with
@@
where $T_{n}$ is a polynomial function of $z$. Note that the sign of $\theta$ may change, depending on the chosen parametrization. Upon setting $T_{n}(z)\sim 1+z$ one obtains the so-called Thomé solutions
@@
By inserting them into Eq. (79) it is possible to obtain a recurrence relation for $\rho_{k},\sigma_{k}$. Note that $\theta_{1},\theta_{2},\theta_{3}$ are in principle different and complex. However, it is possible to relate them to $\alpha,\beta,\gamma,\delta,\eta$ even though this strongly depends on the chosen parametrization.

### v2

@@ VII Large-$r$ Solution of the Full Wave Equation
So far, we have studied the behavior of solutions for fixed frequencies. Since one of our goals is to study the spectrum of a massive field far away from the black hole, we now focus the attention on the study of the solution at spacelike infinity by including the contribution of all frequencies (neglecting the angular part). In the previous sections we have discovered that when $r\rightarrow\infty$, the Fourier modes take the approximate form (49). The large-$r$ form of the scalar field $\psi$ which solves Equation (1) and satisfies the asymptotic condition
@@
is therefore expressed by Equation (4), where, upon defining
@@
The integral is qualitatively different in the $\omega^{2}<\mu^{2}$ case, since the last two terms in Equation (51) become imaginary, leading to
@@
The investigation of these two integrals is a hard task. However, we can rely on analytical approximations and numerical analysis in order to extract the essential features from them. The  starting point is given by studying the behavior in frequencies of the two integrand functions, which we label as
$\Omega_{l\mu}^{I}(t,r,\omega)=\frac{1}{\sqrt{2\pi r(r-2M)}}e^{-i\varphi(t,r,\omega)}\left(1+\sum_{s=1}^{\infty}\frac{B_{s}(l,\omega,\mu)}{r^{s}}\right),$ (54)
@@
Notice that the $l$-dependence is encoded in the coefficients of the series expansion, requiring to add also the angular part (given by spherical harmonics) to study the superposition of several waves. We therefore limit ourselves to the study of the $s$-wave, given by the constraint $B_{s}(l,\omega,\mu)=B_{s}(0,\omega,\mu)$. Moreover, for numerical reasons, we consider only the first five negative powers of $r$, arriving at a simplified version of Equations (54) and (55). Upon setting
@@
we first consider the plots of the real (Figure 7) and imaginary (Figure 8) part of Equation (55).
Figure 7: Plot of Re$\left[\Omega_{0\mu}^{II}(1,r,\omega)\right]$ for $r\in[100,120]$, $\omega\in[-\mu,\mu]$, setting $\mu=0.1$ and $M=10$, up to the $r^{-5}$ term. Along the $\omega$-axis the function has a Gaussian-like behavior (it is possible to prove it by expanding in series Equation (55) for small frequencies), while it diverges along the $r$-axis. The integral, which is, of course, the area under the curve, is clearly divergent in the $r\rightarrow\infty$ limit.
Figure 8: Plot of Im$\left[\Omega_{0\mu}^{II}(1,r,\omega)\right]$ for $r\in[100,120]$, $\omega\in[-\mu,\mu]$, setting $\mu=0.1$ and $M=10$, up to the $r^{-5}$ term. Unlike the real part, the phase shows a different behavior by virtue of its odd parity with respect to $\omega$. The contributions to the integral on the $r$-axis seem to vanish in light of this feature.
It is possible to evaluate numerically the integral given by Equation (53) by expanding $\Omega_{l\mu}^{II}(t,r,\omega)$ around $\omega=0$, since this region gives the largest contribution to it (as shown in Figure 7). We have performed such evaluation up to the sixth power of $\omega$ for the chosen parameters, arriving at the conclusion that such a piece diverges. In order to obtain also an analytical estimate of such a divergence, we consider the change of variable given by
@@
applied to the leading term of Equation (55) and therefore obtain
@@
Even though it is possible to perform more accurate computations (including more terms and/or the negative powers of $r$ of the series expansion) it is clear that the dominant contribution is the one of the exponential term, which causes the divergence of the whole integral in the $r\rightarrow\infty$ limit, as shown by the purely numerical analysis given before. Notice that the field mass term, $\mu$, plays a key role: in the massless case the equivalent of such a piece vanishes (as it happens here in the $\mu\rightarrow 0$ limit). We have to, therefore, face a case where a field not only does not vanish at spacelike infinity, but furthermore it diverges exponentially. As far as we can see, this property is compatible with the findings in Ref. [24], where part (II) of Proposition 2.1 therein shows that the norm of a massive scalar field obeying the wave equation in Schwarzschild space-time is majorized by the product of two functions, one of which is indeed divergent in a suitable limit. This agreement merits explicit mention because the work in Ref.  [24] has exploited advanced methods from the modern theory of hyperbolic equations, whereas we have used the standard techniques of classical mathematical physics.
@@
The case of $\Omega_{l\mu}^{I}(t,r,\omega)$ is strongly different. Upon relying on the same conventions, we show the plots of its real part (the ones for the phase are very similar) for $\omega\in[-10^{3},-\mu]$ in Figure 9 and for $\omega\in[\mu,10^{3}]$ in Figure 10.
@@
On both numerical and analytical grounds, we can say that, while the function $\Omega_{l}^{I}(t,r,\omega)$ diverges, the associated integral (given by Equation (52)) remains finite and eventually vanishes at spacelike infinity. In fact, we have an oscillating term $e^{-i\varphi}$ alleviated by a series of negative powers of $r$.
@@
Interestingly, we find that a massive scalar field contains a term that diverges at spacelike infinity, while this feature is absent in the massless case. In conclusion, the mass plays a crucial role in the asymptotic description of a scalar field in spherical coordinates.
Figure 9: Plot of Re$\left[\Omega_{0\mu}^{I}(1,r,\omega)\right]$ for $r\in[1,1.2]\cdot 10^{3}$, $\omega\in[-10^{3},-\mu]$, setting $\mu=0.1$ and $M=10$, up to the $r^{-5}$ term. For every fixed value of $r$, the field exhibits increasing oscillations as $\omega$ grows (in this specific case for $\omega<-400$). In the low-frequency limit, the field is negligible. Both behaviors suggest that the numerical interval is not divergent.
Figure 10: Plot of Re$\left[\Omega_{0\mu}^{I}(1,r,\omega)\right]$ for $r\in[1,1.2]\cdot 10^{3}$, $\omega\in[\mu,10^{3}]$, setting $\mu=0.1$ and $M=10$, up to the $r^{-5}$ term. The discussion of such a figure follows closely the one of Figure 9.
VIII Application of Section VI to the Zerilli Equation
@@
The systematic algorithm of Section VI can be applied to equations of even greater interest. For this purpose, we here consider the recent analysis in Ref.  [25] of the Zerilli equation [26], which is obtained by perturbing a background Schwarzschild metric with a small parity-even perturbation expressed by a symmetric tensor field of type $(0,2)$. Eventually, the work in Ref. [25] obtains an inhomogeneous radial wave equation, that we express in the form
@@
with $\eta=\frac{(l-1)(l+2)}{2}$. The work in Ref. [25] uses a Green-function method in order to solve the inhomogeneous differential equation (61) by using the well-established technique for obtaining the Green function of the operator on the left-hand side of Equation (61) from solutions of the associated homogeneus equation
@@
It has been our choice to focus on this particular aspect of the general problem since we have already written several sections on the massive scalar case, which is the main topic of our paper.
By repeating the reduction to normal form as we did in Section II, we obtain $U(r,l,\omega)$ in the factorized form
@@
having defined the potential term
@@
At this stage, in close analogy with Equation (41), we may expand Equation (68) as $r\rightarrow\infty$ in the form
@@
The formulae for expressing $\zeta$ and $B_{s}$ in terms of the $A_{k}(l,\omega)$ coincide with the ones in Section VI, whereas the $A_{k}$ are different. Indeed, we find for example that
@@
Hereafter, we plot the real (Figure 11) and imaginary (Figure 12) part of Equation (66) upon considering the parameters
@@
up to the $B_{5}$ coefficient in Equation (71).
@@
Figure 11: The real part of Equation (66) is shown for $r\in[100,300]$, given the parameters (78) and summing over $s$ from 1 to 5 in Equation (71). After achieving high values in the region closest to the event horizon, the field shrinks at $r\approx 140$ and then behaves like a plane wave, as it should be.
@@
Figure 12: The imaginary part of Equation (66) is shown for $r\in[100,300]$, given the parameters (78) and summing over $s$ from 1 to 5 in Equation (71). Its behavior is very similar to the one of the real part, leading to the same conclusions.
Figures 11 and 12 show in a simple way that our findings are in full agreement with the plane-wave behavior as $r\rightarrow\infty$ obtained in Ref. [25] since, after the shrinking of the field at $r\simeq 140$, both real and imaginary part of the solution display constant oscillations through spacelike infinity.
IX Results Obtained and Open Problems
@@
The scalar wave equation in Schwarzschild space-time has been studied intensively over more than half a century [3, 4, 5, 7, 17, 18, 8], until the very recent and highly valuable work in Ref.  [27], but the material in our Sections VI and VII remains original. More precisely, as far as we know, our original contributions deal with ordinary and partial differential equations for black hole space-times and are as follows:
  1. (i)
    For a massive scalar field in Schwarzschild geometry, our Equations (47)–(49) and (108)–(109) yield the full Poincaré asymptotic expansion of Fourier modes that solve the radial wave equation.
  2. (ii)
    Our Equations (52)–(53) yield the integral representation of the massive scalar field, with the resulting numerical analysis discussed therein.
Our formulae provide useful insights into the physical behavior of a massive scalar field in Schwarzschild space-time as the singular points of the radial wave equation are approached. Recognizing the radial equation as nothing but a particular case of the confluent Heun equation makes it possible to rely upon established mathematical techniques without approximations. The algorithm developed in Sections VI and VII allows, in principle, to compute every term of the series at infinity by reaching the desired level of accuracy. Moreover, even if we have shown plots only for specific values of parameters, our formulae can be easily implemented in every symbolic programming language.
@@
For massive scalar fields in Schwarzschild space-time, the literature has mainly focused on late-time behavior [28] of Fourier modes, whereas rigorous results on solutions of the Cauchy problem are along the lines of Ref. [24]. We think that the next step will consist in applying analogous techniques to other spherically symmetric space-times, both in black-hole theory [21] and in cosmological backgrounds [29]. For this purpose, we have eventually studied the homogeneous equation associated with the inhomogeneous Zerilli equation in a Schwarzschild background. Our original Formulas (72)–(77) can be applied to build the Green function, which in turn leads to the solution of the inhomogeneous Zerilli equation as shown in Ref. [25].
More precisely, the following remarks are still in order:
  1. (i)
    The exponential divergence of the massive field at spacelike infinity arises from the integral in Equation (53). As far as we can see, no boundary condition can get rid of it. As  we stressed in Section VII, it is encouraging that our result agrees with a more advanced mathematical analysis as the one performed in Ref. [24], which had been ignored in the physics-oriented literature. Maybe this property means that one has instead to investigate the coupled system consisting of Einstein and Klein–Gordon equations, but this would require a separate paper.
  2. (ii)
    The proof of convergence or divergence of the series in Equation (49) is an open and difficult technical problem because the $B_{s}$ coefficients therein contain rational functions of the parameters and products of such rational functions with square roots of polynomials in the parameters. However, as we say in Appendix E, an even more fundamental open problem is whether one should keep using the Poincaré definition of asymptotic expansion. This topic deserves a separate paper as well. Anyway, for example, the Poincaré asymptotic expansion of the logarithm of the $\Gamma$-function is reliable at large values of the independent variable but is expressed by a divergent series [23]. Thus, good asymptotic estimates at large $r$ in our paper would not be affected by divergence of the expansion.
  3. (iii)
    The cases $\omega^{2}<\mu^{2}$ and $\omega^{2}>\mu^{2}$ studied in Section VII are both physically relevant, and we cannot foresee a viable regularization for the time being. Maybe one has to revert to what we have suggested at the end of item (i) in the present list.
  4. (iv)
    The familiarity with evaluation of $B_{s}$ coefficients gained in Section VI will actually make it easier to complete the analysis of Section VIII. More precisely, the $B_{s}$ coefficients in Equation (71) do not depend on the mass of any field and are simpler in this respect, but on the other hand, they have both real and imaginary parts, as is clear from Equation (77). Their evaluation by computer programs will lead to the evaluation of Fourier modes with the desired accuracy, and in turn, the Green function of the inhomogeneous radial wave equation (61) will be obtained with the same accuracy.
  5. (v)
    We cannot foresee any obstruction to studying wave equations in a Kerr background with our technique. The modern packages will help a lot in evaluating $B_{s}$ coefficients which depend also on angular momentum. Once more, we conclude that a separate paper is necessary to accomplish this task as well.
The authors have equally contributed to conceptualization, methodology, validation, formal analysis. G.E. was the expert in draft preparation, while M.R. was the software expert. All authors have read and agreed to the published version of the manuscript.
This research received no external funding.
No new data were created.
Acknowledgements.
G.E. is grateful to INDAM for membership. G.E. dedicates the paper to his daughter Margherita, and M.R. dedicates this work to his mother Tania.
The authors declare no conflict of interest.
@@ Appendix A The Confluent Heun Equation and Its Solutions
Given a differentiable function $Y$ of the variable $z$, the confluent Heun equation [16] is a second-order ordinary differential equation of the following form:
@@
while $\alpha,\beta,\gamma,\delta,\eta$ are free (in general complex) parameters. A Heun equation has four singular points, while its confluent form has two regular singular points at $z=0$, $z=1$ of $s$-rank $1$ and an irregular singular point at $z=\infty$ of $s$-rank $2$. Indeed, the confluent form is obtained from the standard one by gluing two singularities. In the literature, only power-series solutions in the neighborhood of singular points are known, resulting in three sets of two linearly independent solutions. The ones in the neighborhood of $z=0$ and $z=1$ are called Frobenius solutions, while the ones in the neighborhood of the point at infinity are known as Thomé solutions. Upon denoting with
@@
where $T_{n}$ is a polynomial function of $z$. Note that the sign of $\theta$ may change depending on the chosen parametrization. Upon setting $T_{n}(z)\sim 1+z$ one obtains the so-called Thomé solutions
@@
By inserting them into Equation (79) it is possible to obtain a recurrence relation for $\rho_{k},\sigma_{k}$. Note that $\theta_{1},\theta_{2},\theta_{3}$ are in principle different and complex. However, it is possible to relate them to $\alpha,\beta,\gamma,\delta,\eta$ even though this strongly depends on the chosen parametrization.